OpenStudy (kabbledoo):

Solve −2x2 − 16x − 44 = 0.

OpenStudy (kabbledoo):

Il give a medal plz help

OpenStudy (kabbledoo):

please help

OpenStudy (abdullahm):

What do you need to do?

OpenStudy (aveline):

Start out by dividing by -2

OpenStudy (kabbledoo):

i just need to solve it out im dealing with like imaginary nmbers

OpenStudy (abdullahm):

If we're dealing with imaginary numbers, we should use the quadratic formula. Do you know it?

OpenStudy (kabbledoo):

yea i dont remember it

OpenStudy (abdullahm):

\(\Large\bf x = \frac{-b \pm \sqrt{b^2- 4ac}}{2a}\)

OpenStudy (abdullahm):

Where the equation is ax^2 + bx + c

OpenStudy (aveline):

x^2+8x+22=0 \[\frac{ -8\pm \sqrt{8^{2}-4*1*22} }{ 2*1 }\]

OpenStudy (kabbledoo):

i did it is it x=-4 +/- i(radical)6

OpenStudy (abdullahm):

@Aveline The b term is -16 not 8

OpenStudy (aveline):

\[\frac{ -8\pm 2\sqrt{5} }{ 2 }=\frac{ -4\pm \sqrt{5} }{ 1 }\]

OpenStudy (aveline):

@AbdullahM I simplified by dividing by -2

OpenStudy (abdullahm):

@kabbledoo Solve this: \(\Large\frac{ -(-16)\pm \sqrt{(-16^{2})-4\cdot(-2)\cdot (-44)} }{ 2\cdot(-2) }\)

OpenStudy (kabbledoo):

was i rite?

OpenStudy (abdullahm):

Yes, you are right :)

OpenStudy (abdullahm):

@Aveline I see, my bad. :)

OpenStudy (kabbledoo):

YAY thankss

OpenStudy (abdullahm):

It was my pleasure!